OpenStudy (anonymous):

What is the y-value for f(x) = cos(x) when x = -90°?

OpenStudy (solomonzelman):

There is an identity `cos(-A)=cos(A)`. First, apply this rule.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

I mean, first plug in -90 for x, into cos(x).

OpenStudy (solomonzelman):

and then apply that

OpenStudy (anonymous):

sorry im like also taking notes cuz i dont really get this

OpenStudy (anonymous):

so itll be positive 90 right?

OpenStudy (solomonzelman):

your problem is: \(\large\color{black}{ \displaystyle f(x)= {\rm Cos}(x) }\) plug in -90 for x, \(\large\color{black}{ \displaystyle f(-90)= {\rm Cos}(-90) }\) then since we know that `cos(-a)=cos(a)`, we get: \(\large\color{black}{ \displaystyle f(-90)= {\rm Cos}(90) }\)

OpenStudy (solomonzelman):

Do you know what \(\large\color{black}{ \displaystyle {\rm Cos}(90) }\) is ?

OpenStudy (anonymous):

wait wait hold on

OpenStudy (solomonzelman):

sure, take your time.

OpenStudy (anonymous):

is it -0.44?

OpenStudy (anonymous):

i think im wrong

OpenStudy (solomonzelman):

yes, \(\large\color{black}{ \displaystyle {\rm Cos}(90) }\) is not -0.44

OpenStudy (solomonzelman):

try again please.

OpenStudy (anonymous):

mmmm 51

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

no idk wats cos (90) ....im so sorry i really am trying :(

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm Cos}(90)=1 }\)

OpenStudy (anonymous):

why

OpenStudy (solomonzelman):

that is a good question. \(\large\color{black}{ \displaystyle {\rm Cos}(90)={\rm Cos}(45+45) ={\rm Cos}(45) {\rm Cos}(45) -{\rm Sin}(45) {\rm Sin}(45) }\) \(\large\color{slate}{ \\[0.5em] }\) `sine = opposite / hypotenuse` `cosine = adjacent / hypotenuse`\(\large\color{slate}{ \\[0.8em] }\) since opposite and adjacent sides are formed \(\large\color{slate}{ \\[0.8em] }\) by the same angle, they are (themselves) the same.\(\large\color{slate}{ \\[0.8em] }\) Saying that `opposite=adjacent`\(\large\color{slate}{ \\[0.8em] }\) And if `opposite=adjacent`, \(\large\color{slate}{ \\[0.8em] }\) then `opposite / hypotenuse = adjacent / hypotenuse` \(\large\color{slate}{ \\[0.8em] }\) which is essentially telling us that `sin(45)=cos(45)`, \(\large\color{slate}{ \\[0.8em] }\) because `sine = opposite / hypotenuse` and `cosine = adjacent / hypotenuse` Then continuing on our math, \(\large\color{black}{ \displaystyle {\rm Cos}(90)={\rm Cos}(45+45) ={\rm Cos}(45) {\rm Cos}(45) -{\rm Sin}(45) {\rm Sin}(45) }\) if `sin(45)=cos(45)` , then `sin(45)sin(45)=cos(45)cos(45)` and therefore, \(\large\color{black}{ \displaystyle {\rm Cos}(45) {\rm Cos}(45) -{\rm Sin}(45) {\rm Sin}(45)\color{blue}{=0} }\)